### v1

@@ Abstract
We present the discovery of NGC253-SNFC-dw1, a new satellite galaxy in the remote stellar halo of the Sculptor Group spiral, NGC 253. The system was revealed using deep resolved star photometry obtained as part of the Subaru Near-Field Cosmology Survey that uses the Hyper Suprime-Cam on the Subaru Telescope. Although rather luminous () and massive (), the system is one of the most diffuse satellites yet known, with a half-light radius of kpc and an average surface brightness of mag arcmin within the . The colour-magnitude diagram shows a dominant old ( Gyr) and metal-poor ( dex) stellar population, as well as several candidate thermally-pulsing asymptotic giant branch stars. The distribution of red giant branch stars is asymmetrical and displays two elongated tidal extensions pointing towards NGC 253, suggestive of a highly disrupted system being observed at apocenter. NGC253-SNFC-dw1 has a size comparable to that of the puzzling Local Group dwarfs Andromeda XIX and Antlia 2 but is two magnitudes brighter. While unambiguous evidence of tidal disruption in these systems has not yet been demonstrated, the morphology of NGC253-SNFC-dw1 clearly shows that this is a natural path to produce such diffuse and extended galaxies. The surprising discovery of this system in a previously well-searched region of the sky emphasizes the importance of surface brightness limiting depth in satellite searches.
@@ 1 Introduction
The CDM cosmological model predicts that galaxies form hierarchically; large galaxies like the Milky Way originate from small over-densities in the primordial matter distribution and grow via the agglomeration of numerous smaller building blocks, some of which survive later merging and represent the present-day dwarf satellites (e.g. Bullock & Johnston, 2005; Pillepich et al., 2014). The numbers and properties of satellite galaxies around massive hosts has emerged as a powerful way to test the CDM paradigm on small scales as well as the physics of galaxy formation (e.g. Bullock & Boylan-Kolchin, 2017; Sales et al., 2022). As a result, there has been much interest in characterising the satellite system of the Milky Way and its nearest neighbours (e.g. Homma et al., 2019; Carlsten et al., 2022; Doliva-Dolinsky et al., 2023). Although the sample size beyond the Local Group (LG) is still limited, such observations are critical to benchmark results from the Milky Way and M31. Recent wide-field imaging and/or survey campaigns are having remarkable success in uncovering new satellites around Local Volume galaxies (e.g. Chiboucas et al., 2013; Crnojević et al., 2016; Okamoto et al., 2019; Mutlu-Pakdil et al., 2022) but these studies rapidly lose sensitivity at surface brightnesses mag arcsec, raising questions about what they might be missing.
@@
In recent years, attention has focused on the faint satellite system of NGC 253. The PISCeS survey used Magellan/Megacam resolved star photometry to discover five new dwarf satellites of NGC 253 (Sand et al., 2014; Toloba et al., 2016; Mutlu-Pakdil et al., 2022), with two of them, Scl-MM-dw2 (hereafter MM-dw2) and Scl-MM-dw3, independently discovered as NGC253-dw2 by Romanowsky et al. (2016) and Donatiello II by Martínez-Delgado et al. (2021), respectively. Martínez-Delgado et al. (2021) proposed two further dwarf galaxy candidates from Dark Energy Survey data and Carlsten et al. (2022) suggested two others from DECaLS data, both from integrated light analyses. Mutlu-Pakdil et al. (2024) confirm three of these latter systems with Hubble Space Telescope photometry and undertake a comprehensive study of the NGC 253 satellite system as a whole, finding it to be deficient in luminous satellites compared systems of similar stellar mass.
@@ 3.2 Distance
We derive the distance to SNFC-dw1 using the TRGB method (e.g. Lee et al., 1993), following the procedure described in Okamoto et al. (2019). Stars of that lie within 9 arcmin radius are used to derive the -band luminosity function (LF). Using the LF of point sources in the reference field, we correct for contamination from the NGC 253 halo population and from foreground/background objects. Then, we apply a Sobel filter to detect a sharp transition at . We estimate the TRGB color as and apply equation (1) of Okamoto et al. (2019) to calculate as . Therefore, the distance modulus to SNFC-dw1 is , corresponding to D Mpc. To be able to directly compare, we also estimate the TRGB distance of NGC 253 using our dataset. In this case, we use stars lying in the elliptical annulus covering and correct for contamination of foreground/background objects using the same reference field but scaled up to cover the same area within the elliptical annulus. We find , or D Mpc, which is in excellent agreement with Radburn-Smith et al. (2011). This places SNFC-dw1 at a 3D radius of kpc on the far side of NGC 253.
@@ 3.3 Luminosity, Structure and Metallicity
Figure 3 shows the metallicity distribution function (MDF) of SNFC-dw1 derived from the completeness-corrected color distribution of bright RGB stars. We use RGB stars of to reduce the uncertainties and to keep the wide spacing between isochrones in the color. RGB stars are selected within the solid circle of the right panel of Figure 1 and within the solid polygon of Figure 2. The metallicity is derived from the linear interpolation of isochrones of to with a fixed age of 10 Gyr and extrapolation with the spline function up to in the same manner as done in Okamoto et al. (2023). The completeness is corrected as before according to the color and magnitude of each star. The same estimation is done for the reference field in order to derive the contaminant contribution to the MDF arising from the NGC 253 halo and background/foreground objects. The contaminant-corrected MDF is shown as the shaded histogram. The uncertainty is estimated by performing a Monte Carlo simulation with N for each RGB star. A star is randomly re-sampled from a Gaussian distribution with a width equal to the photometric error, and the metallicity is redetermined. The standard deviation of the resulting metallicity distribution is adopted as the uncertainty for each RGB star. The estimated uncertainties increase from 0.03 to 0.14 dex with decreasing metallicity from to due to the narrow range in RGB color between metal-poor isochrones. The weighted mean metallicity corrected for this effect is shown as the error-weighted mean value dex. The overall appearance of the MDF is well-described by a single Gaussian shown as the solid line in Figure 3. The sudden drop on the metal-rich side () of the MDF likely reflects an over-subtraction due to small number statistics in this metallicity range, in which the NGC 253 halo dominates the contaminant signal.
@@
Although the extended regions of SNFC-dw1 are very irregular, the core structure seen in the lower-middle panel of Figure 1 is well-defined and so we proceed to estimate its structural properties. We compute the values using the density-weighted zeroth, first and second moments of the metal-poor RGB spatial distribution and list them in Table 1. The estimated shape shown as a dashed line in Figure 1 seems to be more circular than the visual impression of the density map, probably due to the existence of extended components toward the north. We also fit the standard Sérsic profile via least-squares minimization to the completeness-corrected metal-poor RGB star counts. The radial profile is constructed by calculating the average number density of RGB stars in a series of elliptical annuli, defined using the derived structural parameters. We take the Poisson uncertainties in the RGB counts as the uncertainty. Initially, the contaminant contribution is visually estimated as the average number density at large radius () and we fit the profile and estimate the half-light radius (). We then iterate on the estimation of the contaminant level, this time using the average number density between 4 and . Finally, the profile is refit with this contamination level. The resulting kpc is very large for a dwarf galaxy and, in combination with the relatively low luminosity, yields an extremely faint average surface brightness of mag arcsec within .
@@ 4 Discussion and Summary
Using data from the SNFC survey being conducted with the Subaru Telescope, we have uncovered a new dwarf satellite of the nearby spiral NGC 253. The system lies at a projected separation of 83.4 kpc from the center of NGC 253. We have used the TRGB method to derive consistent distances to SNFC-dw1 and NGC 253, finding them to lie at Mpc and Mpc, respectively. This places SNFC-dw1 at a 3D radius of 146 kpc on the far side of the NGC 253 halo and well within its estimated virial radius of 330 kpc (Mutlu-Pakdil et al., 2021), confirming it is a bound satellite.
@@
We find SNFC-dw1 to be dominated by an old metal-poor population; additionally, there is tentative evidence for an associated TP-AGB population. As such stars have ages of typically Gyr (Marigo et al., 2017), their presence suggests that the system was star-forming until relatively recently. Although SNFC-dw1 has a comparable luminosity, stellar mass and metallicity to bright classical Galactic dSphs, it is extremely diffuse ( mag arcsec) – a direct consequence of its abnormally large size with kpc. Indeed, star counts can be traced over more than 10 kpc and reveal a very asymmetric structure in the peripheral regions. Notably, SNFC-dw1 has a size comparable to some of the most diffuse galaxies currently known, including the puzzling Local Group systems Crater 2 ( kpc), Antlia 2 ( kpc) and Andromeda XIX ( kpc) (Torrealba et al., 2016; Ji et al., 2021; Savino et al., 2022).
@@
Figure 4 compares SNFC-dw1 to other dwarf galaxies in terms of scaling relationships. These plots include LG dSphs (McConnachie, 2012, updated in January 2021), Ultra-diffuse galaxies (UDGs) in the Coma Cluster (van Dokkum et al., 2015), some of the known NGC 253 satellites (Toloba et al., 2016; Martínez-Delgado et al., 2021; Mutlu-Pakdil et al., 2022) and the particularly diffuse dSphs (open squares) Antlia 2 (Torrealba et al., 2019; Ji et al., 2021; Vivas et al., 2022), Andromeda XIX (Collins et al., 2020), the M81 satellite F8D1 (Žemaitis et al., 2023), the Centauras A satellite CenA-MM-dw1 (Crnojević et al., 2019), and the recently-discovered NGC 55-dw1 (McNanna et al., 2024). The upper panel of Figure 4 shows the systems on the size-luminosity plane. Only six systems have a similarly large size ( kpc) to SNFC-dw1 – Andromeda XIX and Antlia 2 and NGC 55-dw1, which are magnitudes fainter, and the Sagittarius (Sgr) dSph, CenA-MM-dw3 and F8D1, which are magnitudes brighter. While both Sgr, CenA-MM-dw3 and F8D1 show direct evidence for tidal disruption (Crnojević et al., 2016; Majewski et al., 2003; Žemaitis et al., 2023), the evidence in Antlia 2 and Andromeda XIX has thus far only been circumstantial (e.g. Belokurov et al., 2006; Torrealba et al., 2019; Collins et al., 2020; Ji et al., 2021). The situation is also not clear for NGC 55-dw1 which, in addition to lacking clear tidal features, does not yet have a directly measured distance (McNanna et al., 2024). Amongst the NGC 253 satellites, MM-dw2 is most similar to SNFC-dw1 and it exhibits a highly elongated shape (see Figure 1). However, the properties of MM-dw2 estimated from two independent studies are rather different – and kpc using Magellan/Megacam (Toloba et al., 2016) and and kpc using Subaru/Suprime-Cam (Romanowsky et al., 2016). We plot both values connected with a purple solid line in the upper panel of Figure 4. The prominent tidal extensions emanating from SNFC-dw1 provide unambiguous evidence of tidal disruption, demonstrating that this is a natural pathway to produce such extremely diffuse and extended galaxies. This conclusion is further supported by integrated light analyses of two more distant dwarfs with similarly large sizes – NGC4449B associated with the dwarf starburst galaxy NGC 4449 (Rich et al., 2012; Martínez-Delgado et al., 2012) and HCC-087 in the Hydra I galaxy cluster, (Koch et al., 2012), both of which are also known to exhibit S-shaped tidal features.
@@
It is worth noting that SNFC-dw1 lies in an area of the sky that has been the target of multiple dwarf galaxy searches in recent years, some of which have used integrated light (Martínez-Delgado et al., 2021; Mutlu-Pakdil et al., 2022) while others have used resolved stars (Carlsten et al., 2022). Although it is the second most luminous of the six satellites projected within 150 kpc of NGC 253 (Mutlu-Pakdil et al., 2024), and comparable to the brightest classical dSph of the Milky Way and Andromeda, SNFC-dw1 has remained unearthed until now due to its extremely diffuse nature. This emphasizes the importance of surface brightness limiting depth in satellite searches, suggesting that even bright-end incompleteness can be an issue at the typical depths of most present-day studies. Fortunately, with the recent launch of ESA’s Euclid satellite, the start of the Legacy Survey of Space and Time with the Vera C. Rubin Observatory in 2025 and the Nancy Grace Roman Space Telescope launch in 2027, the era of very sensitive ( mag arcsec) satellite searches around large numbers of Local Volume galaxies will soon be upon us.

### v2

@@ Abstract
We present the discovery of NGC253-SNFC-dw1, a new satellite galaxy in the remote stellar halo of the Sculptor Group spiral, NGC 253. The system was revealed using deep resolved star photometry obtained as part of the Subaru Near-Field Cosmology Survey that uses the Hyper Suprime-Cam on the Subaru Telescope. Although rather luminous () and massive (), the system is one of the most diffuse satellites yet known, with a half-light radius of kpc and an average surface brightness of mag arcmin-2 within the . The colour-magnitude diagram shows a dominant old ( Gyr) and metal-poor ( dex) stellar population, as well as several candidate thermally-pulsing asymptotic giant branch stars. The distribution of red giant branch stars is asymmetrical and displays two elongated tidal extensions pointing towards NGC 253, suggestive of a highly disrupted system being observed at apocenter. NGC253-SNFC-dw1 has a size comparable to that of the puzzling Local Group dwarfs Andromeda XIX and Antlia 2 but is two magnitudes brighter. While unambiguous evidence of tidal disruption in these systems has not yet been demonstrated, the morphology of NGC253-SNFC-dw1 clearly shows that this is a natural path to produce such diffuse and extended galaxies. The surprising discovery of this system in a previously well-searched region of the sky emphasizes the importance of surface brightness limiting depth in satellite searches.
@@ 1 Introduction
The CDM cosmological model predicts that galaxies form hierarchically; large galaxies like the Milky Way originate from small over-densities in the primordial matter distribution and grow via the agglomeration of numerous smaller building blocks, some of which survive later merging and represent the present-day dwarf satellites (e.g. Bullock & Johnston, 2005; Pillepich et al., 2014). The numbers and properties of satellite galaxies around massive hosts has emerged as a powerful way to test the CDM paradigm on small scales as well as the physics of galaxy formation (e.g. Bullock & Boylan-Kolchin, 2017; Sales et al., 2022). As a result, there has been much interest in characterising the satellite system of the Milky Way and its nearest neighbours (e.g. Homma et al., 2019; Carlsten et al., 2022; Doliva-Dolinsky et al., 2023). Although the sample size beyond the Local Group (LG) is still limited, such observations are critical to benchmark results from the Milky Way and M31. Recent wide-field imaging and/or survey campaigns are having remarkable success in uncovering new satellites around Local Volume galaxies (e.g. Chiboucas et al., 2013; Crnojević et al., 2016; Okamoto et al., 2019; Mutlu-Pakdil et al., 2022) but these studies rapidly lose sensitivity at surface brightnesses mag arcsec-2, raising questions about what they might be missing.
@@
In recent years, attention has focused on the faint satellite system of NGC 253. The PISCeS survey used Magellan/Megacam resolved star photometry to discover five new dwarf satellites of NGC 253 (Sand et al., 2014; Toloba et al., 2016; Mutlu-Pakdil et al., 2022), with two of them, Scl-MM-dw2 (hereafter MM-dw2) and Scl-MM-dw3, independently discovered as NGC253-dw2 by Romanowsky et al. (2016) and Donatiello II by Martínez-Delgado et al. (2021), respectively. Martínez-Delgado et al. (2021) proposed two further dwarf galaxy candidates from Dark Energy Survey data and Carlsten et al. (2022) suggested another one from DECaLS data, both from integrated light analyses. Mutlu-Pakdil et al. (2024) confirm these latter systems with Hubble Space Telescope photometry and undertake a comprehensive study of the NGC 253 satellite system as a whole, finding it to be deficient in luminous satellites compared systems of similar stellar mass.
@@ 3.2 Distance
We derive the distance to SNFC-dw1 using the TRGB method (e.g. Lee et al., 1993), following the procedure described in Okamoto et al. (2019). Stars of that lie within 9 arcmin radius are used to derive the -band luminosity function (LF). Using the LF of point sources in the reference field, we correct for contamination from the NGC 253 halo population and from foreground/background objects. Then, we apply a Sobel filter to detect a sharp transition at . We estimate the TRGB color as and apply equation (1) of Okamoto et al. (2019) to calculate as . Therefore, the distance modulus to SNFC-dw1 is , corresponding to D Mpc. To be able to directly compare, we also estimate the TRGB distance of NGC 253 using our dataset. In this case, we use stars lying in the elliptical annulus covering and correct for contamination of foreground/background objects using the same reference field but scaled up to cover the same area within the elliptical annulus. We find , or D Mpc, which is in excellent agreement with Radburn-Smith et al. (2011). This places SNFC-dw1 at a 3D radius of kpc on the far side of NGC 253 but the uncertainties are sufficiently large to make it very tentative.
@@ 3.3 Luminosity, Structure and Metallicity
Figure 3 shows the metallicity distribution function (MDF) of SNFC-dw1 derived from the completeness-corrected color distribution of bright RGB stars. We use RGB stars of to reduce the uncertainties and to keep the wide spacing between isochrones in the color. RGB stars are selected within the solid circle of the right panel of Figure 1 and within the solid polygon of Figure 2. The metallicity is derived from the linear interpolation of isochrones of to with a fixed age of 10 Gyr and extrapolation with the spline function up to in the same manner as done in Okamoto et al. (2023). The completeness is corrected as before according to the color and magnitude of each star. The same estimation is done for the reference field in order to derive the contaminant contribution to the MDF arising from the NGC 253 halo and background/foreground objects. The contaminant-corrected MDF is shown as the shaded histogram. The uncertainty is estimated by performing a Monte Carlo simulation with N for each RGB star. A star is randomly re-sampled from a Gaussian distribution with a width equal to the photometric error, and the metallicity is redetermined. The standard deviation of the resulting metallicity distribution is adopted as the uncertainty for each RGB star. The estimated uncertainties increase from 0.03 to 0.14 dex with decreasing metallicity from to due to the narrow range in RGB color between metal-poor isochrones. The weighted mean metallicity corrected for this effect is shown as the error-weighted mean value dex. The overall appearance of the MDF is well-described by a single Gaussian shown as the solid line in Figure 3. The metallicity dispersion is significant () compared to the uncertainty (0.04) on the fit. The sudden drop on the metal-rich side () of the MDF likely reflects an over-subtraction due to small number statistics in this metallicity range, in which the NGC 253 halo dominates the contaminant signal.
@@
Although the extended regions of SNFC-dw1 are very irregular, the core structure seen in the lower-middle panel of Figure 1 is well-defined and so we proceed to estimate its structural properties. We compute the values using the density-weighted zeroth, first and second moments of the metal-poor RGB spatial distribution and list them in Table 1. The estimated shape shown as a dashed line in Figure 1 seems to be more circular than the visual impression of the density map, probably due to the existence of extended components toward the north. We also fit the standard Sérsic profile via least-squares minimization to the completeness-corrected metal-poor RGB star counts. The radial profile is constructed by calculating the average number density of RGB stars in a series of elliptical annuli, defined using the derived structural parameters. We take the Poisson uncertainties in the RGB counts as the uncertainty. Initially, the contaminant contribution is visually estimated as the average number density at large radius () and we fit the profile and estimate the half-light radius (). We then iterate on the estimation of the contaminant level, this time using the average number density between 4 and . Finally, the profile is refit with this contamination level. The resulting kpc is very large for a dwarf galaxy and, in combination with the relatively low luminosity, yields an extremely faint average surface brightness of mag arcsec-2 within .
@@ 4 Discussion and Summary
Using data from the SNFC survey being conducted with the Subaru Telescope, we have uncovered a new dwarf satellite of the nearby spiral NGC 253. The system lies at a projected separation of 83.4 kpc from the center of NGC 253. We have used the TRGB method to derive consistent distances to SNFC-dw1 and NGC 253, finding them to lie at Mpc and Mpc, respectively. This places SNFC-dw1 well within its estimated virial radius of 330 kpc (Mutlu-Pakdil et al., 2021), confirming it is a bound satellite.
@@
We find SNFC-dw1 to be dominated by an old metal-poor population; additionally, there is tentative evidence for an associated TP-AGB population. As such stars have ages of typically Gyr (Marigo et al., 2017), their presence suggests that the system was star-forming until relatively recently. Although SNFC-dw1 has a comparable luminosity, stellar mass, and metallicity to bright classical Galactic dSphs, it is extremely diffuse ( mag arcsec-2) – a direct consequence of its abnormally large size with kpc. Indeed, star counts can be traced over more than 10 kpc and reveal a very asymmetric structure in the peripheral regions. Notably, SNFC-dw1 has a size comparable to some of the most diffuse galaxies currently known, including the puzzling Local Group systems Crater 2 ( kpc), Antlia 2 ( kpc) and Andromeda XIX ( kpc) (Torrealba et al., 2016; Ji et al., 2021; Savino et al., 2022).
@@
Figure 4 compares SNFC-dw1 to other dwarf galaxies in terms of scaling relationships. These plots include LG dSphs (McConnachie, 2012, updated in January 2021), Ultra-diffuse galaxies (UDGs) in the Coma Cluster (van Dokkum et al., 2015), some of the known NGC 253 satellites (Toloba et al., 2016; Martínez-Delgado et al., 2021; Mutlu-Pakdil et al., 2022) and the particularly diffuse dSphs (open squares) Antlia 2 (Torrealba et al., 2019; Ji et al., 2021; Vivas et al., 2022), Andromeda XIX (Collins et al., 2020), the M81 satellite F8D1 (Žemaitis et al., 2023), the Centauras A satellite CenA-MM-dw1 (Crnojević et al., 2019), and the recently-discovered NGC 55-dw1 (McNanna et al., 2024). The upper panel of Figure 4 shows the systems on the size-luminosity plane. Only six systems have a similarly large size ( kpc) to SNFC-dw1 – Andromeda XIX and Antlia 2 and NGC 55-dw1, which are magnitudes fainter, and the Sagittarius (Sgr) dSph, CenA-MM-dw3 and F8D1, which are magnitudes brighter. While both Sgr, CenA-MM-dw3 and F8D1 show direct evidence for tidal disruption (Crnojević et al., 2016; Majewski et al., 2003; Žemaitis et al., 2023), the evidence in Antlia 2 and Andromeda XIX has thus far only been circumstantial (e.g. Belokurov et al., 2006; Torrealba et al., 2019; Collins et al., 2020; Ji et al., 2021). The situation is also not clear for NGC 55-dw1 which, in addition to lacking clear tidal features, does not yet have a directly measured distance (McNanna et al., 2024). Amongst the NGC 253 satellites, MM-dw2 is most similar to SNFC-dw1 and it exhibits a highly elongated shape (see Figure 1). However, the properties of MM-dw2 estimated from two independent studies are rather different – and kpc using Magellan/Megacam (Toloba et al., 2016) and and kpc using Subaru/Suprime-Cam (Romanowsky et al., 2016). We plot both values connected with a purple solid line in the upper panel of Figure 4. The prominent tidal extensions emanating from SNFC-dw1 provide unambiguous evidence of tidal disruption, demonstrating that this is a natural pathway to produce such extremely diffuse and extended galaxies. This conclusion is further supported by integrated light analyses of two more distant dwarfs with similarly large sizes – NGC4449B associated with the dwarf starburst galaxy NGC 4449 (Rich et al., 2012; Martínez-Delgado et al., 2012) and HCC-087 in the Hydra I galaxy cluster (Koch et al., 2012), both of which are also known to exhibit S-shaped tidal features.
@@
It is worth noting that SNFC-dw1 lies in an area of the sky that has been the target of multiple dwarf galaxy searches in recent years, some of which have used integrated light (Martínez-Delgado et al., 2021; Mutlu-Pakdil et al., 2022) while others have used resolved stars (Carlsten et al., 2022). Although it is the second most luminous of the six satellites projected within 150 kpc of NGC 253 (Mutlu-Pakdil et al., 2024), and comparable to the brightest classical dSph of the Milky Way and Andromeda, SNFC-dw1 has remained unearthed until now due to its extremely diffuse nature. This emphasizes the importance of surface brightness limiting depth in satellite searches, suggesting that even bright-end incompleteness can be an issue at the typical depths of most present-day studies. Fortunately, with the recent launch of ESA’s Euclid satellite, the start of the Legacy Survey of Space and Time with the Vera C. Rubin Observatory in 2025 and the Nancy Grace Roman Space Telescope launch in 2027, the era of very sensitive ( mag arcsec-2) satellite searches around large numbers of Local Volume galaxies will soon be upon us.
5 Acknowledgments
We thank Burçin Mutlu-Pakdil, David Sand and the anonymous referee for helpful comments.
SO acknowledges support in part from JSPS Grant-in-Aid for Scientific Research (18H05875, 20K04031, 20H05855). AMNF and RZ are grateful for support from the UK STFC via grants ST/Y001281/1 and ST/V000594/1. JK acknowledges support from NSF through grants AST-1812847 and AST-2006600. MGL was supported by the National Research Foundation grant funded by the Korean Government (NRF-2019R1A2C2084019).
We are grateful to all the staff at Subaru Telescope and the HSC team. This research is based on data collected at the Subaru Telescope, operated by the National Astronomical Observatory of Japan. We are honored and grateful for the opportunity to observe the Universe from Maunakea, which has cultural, historical, and natural significance in Hawaii.
This paper makes use of software developed for Vera C. Rubin Observatory. We thank the Rubin Observatory for making their code available as free software at http://pipelines.lsst.io/. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation.
For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.